Supplemental Operations Report 03Dec2022

[title Supplemental Operations Report – December 3rd]

Name of person filing report: Shannon Rupert

Reason for Report: Routine

Non-nominal systems: Generator

Action taken for non-nominal systems: Propane regulator was checked and pressure increased from 9 to 11. The generator seems to be running better, but there is still some sort of miscommunication between the generator and power system. Still working on it.

Generator: Running from about 8 pm to 8 am.

ScienceDome Dual Split: Heat on between 8 pm and 8 am at 67degrees F.

Solar— Nominal.
Solar— SOC Last 24 hours:
Max 100%
Min 52%
Avg 70.0% (under heavy cloud cover)
Propane Reading, station tank – filled today %
Propane Reading, director tank— filled today %
Propane Reading, intern tank— filled today %
Propane Reading, generator— filled today %

Ethanol Free Gasoline – 0 gallons

Water (static tank) – 175 gallons

Water in GreenHab – Not noted gallons

Water in ScienceDome: 0 gallons

Water (Outpost tank) – 175 gallons

Hab toilet tank emptied: check regular operations reports

Sojourner rover used: yes

Hours: not noted

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: We disconnected the heat sensor from the motor in Oppy. Oppy has had a faulty sensor since it was new (this is a factory problem for some Polaris Ranger EVs). As a result, Oppy has been underused. The only way to fix the problem was to replace the motor, which would cost $2800. Instead we disconnected the sensor, then tested the rover by driving it as fast and as safely as possible to the cattle guard at the beginning of Cow Dung Road and back. We checked the motor temperature by touch several times and each time it was only as hot as Soji. The only downside to this repair is that if the motor does actually overheat in the future, the motor will fail. Then we will have to replace it, but if this repair holds we will have a fully functioning rover. It currently has less than half the hours of every other rover, and now it can be used more.

ATV’s Used: (Honda, 350.1, 350.2, 300): Honda

Reason for use: charging battery

Oil Added? no

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: Currently waiting to be picked up for serving. After I started and moved the Honda today, it got stuck in 2nd gear. This means that none of the ATVs are currently operational.

HabCar used and why, where? Yes, to town and to the dump

CrewCar used and why, where? Yes, to town and to Grand Junction

Luna used and why, where? Yes, to town and to Loa

General notes and comments: Tabs for the year were purchased for the 2018 rovers and the HabCar. That means all vehicles are registered until next year.

Summary of internet: Nominal. HudgesNet is currently off, and we plan to reroute it for crew use and keep Starlink at the Outpost except when crews are out of sim. The new iMac has been installed at the HabCom Station on the lower deck, however we will be moving it back to the Outpost and relocating the desk back to the SciDome when this crew is completed. We need to reimagine the HabCom station because where we put it, it was just getting used as a workbench.

Summary of suits and radios: See crew operations reports. The next upgrade to the repeater system will take place when Crew 271 is on station.

Campus wide inspection, if action taken, what and why? There are great advantages to having two people working on campus. Today we hitched the trash trailer up to the HabCar and delivered years of accumulated waste to the dump. The trailer still has the one tire with the slow leak but otherwise worked well. We added air to all tires so that each was at 30 psi.

Summary of general operations: Started work on moving the conduit.

Summary of Hab operations: Nothing to report

Summary of Outpost operations: Nothing to report

Summary of GreenHab operations: Plants were fertilized. Tomato cages were installed. Timer was installed for four hours of supplemental light from 10 pm – 2 am.

Summary of ScienceDome operations: Nothing to report. Duel split turned on to 67 degrees F overnight, same timing as the generator.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns, supplies needed and requests: n/a

Copyright © The Mars Society. All rights reserved. | Main Site